Hello out there,I need some help with some problems,the first one is,Factor.Write prime if the expression cannot be factored:49y^2-25.The next one is,Factor.Write prime if the expression cannot be factored:4x^2-24x+36.The next one is,Factor.Write prime if the expression cannot be factored:16e^2-4.And the last one is,Factor.Write prime if the expression cannot be factored:4k^2+20k+25.Thanks to whoever can help.
Answer by stanbon(57347)

first one is,Factor.Write prime if the expression cannot be factored:
49y^2-25 = (7y-5)(7y+5)
--------------
The next one is,Factor.Write prime if the expression cannot be factored:
4x^2-24x+36 = 4(x^2-6x+9) = 4(x-3)(x-3)
-----------------
The next one is,Factor.Write prime if the expression cannot be factored:
16e^2-4 = (4e-2)(4e+2)
-----------------
And the last one is,Factor.Write prime if the expression cannot be factored:
4k^2+20k+25 = (2k+5)(2k+5)
===============
Cheers,
Stan H.